OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (0, -8) and a directrix at y = 8.

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

this is real real easy if you know what it looks like do you know that?

OpenStudy (anonymous):

\[y=a(x-h)^2+k\]

OpenStudy (anonymous):



OpenStudy (misty1212):

i was asking if you knew what the parabola looked like, not the form of it, but rather the picture of it

OpenStudy (misty1212):

@satellite73 drew the directrix and the focus do you know how the parabola looks? i.e. opens up, down, left or right? also the vertex?

OpenStudy (misty1212):

if so, we can answer this question in like 2 seconds if not, it is almost impossible

OpenStudy (anonymous):

vertex would be (0,0) right?

OpenStudy (misty1212):

yes

OpenStudy (misty1212):

does it open up or down?

OpenStudy (anonymous):

up?

OpenStudy (misty1212):

hmm no

OpenStudy (misty1212):



OpenStudy (anonymous):

oh sorry down, i mixed up the a>0 and a<0

OpenStudy (misty1212):

it opens down because the focus is below the directrix

OpenStudy (misty1212):

that means it is going to look like \[-4p(y-k)=(x-h)^2\] but since \((h,k)=(0,0)\) it is just \[-4py=x^2\]

OpenStudy (misty1212):

\(p\) is the distance between the focus and the vertex, which is evidently 8

OpenStudy (misty1212):

so \[-4\times 8y=x^2\] or \[-32y=x^2\] and we are done

OpenStudy (anonymous):

the only answer i have that is similar to that is y^2= -32x

OpenStudy (misty1212):

that is why it is important to know what it looks like first, so you know the right form to put it in

OpenStudy (anonymous):

is it the same?

OpenStudy (misty1212):

is one choice \[y=-\frac{1}{32}x^2\]

OpenStudy (anonymous):

yes

OpenStudy (misty1212):

no they are not the same, but \[-32y=x^2\] is the same as \[y=-\frac{1}{32}x^2\]

OpenStudy (anonymous):

okay thank you very much!

OpenStudy (misty1212):

i.e. divide both sides by \(-32\)

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (anonymous):

is there a way you could help me on a few other questions?

OpenStudy (misty1212):

if i know how go ahead and post one

OpenStudy (misty1212):

you can post it here if you like